# Wireless 6G Connectivity for Massive Number of Devices and Critical Services

### v2

@@ II-C2 Covariance-Based Decoding
The covariance-based approach arises from the observation that the received signal at each antenna are independent and identically distributed realizations of a multivariate complex Gaussian distribution with zero mean and covariance matrix $\mathbf{\Sigma}=\sum_{k=1}^{N}\gamma_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}+\sigma_{Z}^{2}\mathbf{I}$, where the superscript $H$ stands for conjugate transposition. The likelihood function of the received signal $\mathbf{Y}$ given $\mathbf{\gamma}$ can then be computed as
$p(\mathbf{Y}|\mathbf{\gamma})$ $=\prod_{m=1}^{M}\frac{1}{|\pi\mathbf{\Sigma}|}\exp\left(-\mathbf{y}_{m}^{H}\mathbf{\Sigma}^{-1}\mathbf{y}_{m}\right)$ (13)
@@
Although maximizing the likelihood leads to a non-convex problem, it is possible to obtain good solutions using a coordinate descent algorithm. The resulting set of active users can then be recovered from $\mathbf{\gamma}$. By analyzing the coordinate descent, it can be shown that the computational complexity is $O(TNL^{2})$, where $T$ is the number of iterations [19]. In practice, the covariance-based method tends to have longer runtime than AMP for small values of $M$, but is more efficient for larger $M$ [23].
Compared to AMP, performance analysis of the covariance-based approach is more involved, but have been characterized in [23] for the asymptotic case where the number of antennas goes to infinity while $N$, $K$, and $L$ remain finite. In this regime, it is possible to numerically obtain the detection error. This analysis reveals that the algorithm exhibits a phase transition, where the decoding error vanishes for a specific region of $N$, $K$, and $L$. As a key advantage as compared to AMP, the covariance-based approach can work well in the regime where the number of users $K$ is in the order of the square of the preamble length $L$. This is due to the fact that the covariance matrix is of size $O(L^{2})$. However, such advantage comes at a cost of not being able to estimate the channels. Finally, we remark that while the discussion in this session has been confined to the single-cell case, both the algorithmic developments and the analyses of AMP and the covariance method can be extended to the multi-cell or the cell-free massive multiple-input multiple-output (MIMO) scenarios [24, 25, 26, 27].
@@ II-D Massive Downlink Connectivity
Figure 4: A general massive downlink channel. Only a small random subset comprising $K$ messages are informative ($W_{n}\in\mathcal{M}$), while the remaining $K-N$ messages are empty ($W_{n}=\varnothing$).
A general representation of the massive downlink channel is depicted in Fig. 4, where the base station aims to jointly send $N$ messages $W_{1},\ldots,W_{N}$ to users $1,\ldots,N$, respectively, where $W_{n}\in\mathcal{M}\cup\{\varnothing\}$ for $n=1,\ldots,N$. Each message $W_{n}$ ($n=1,\ldots,N$) can either be informative, drawn from the message set $\mathcal{M}$, or empty, denoted by $W_{n}=\varnothing$, indicating that the transmitter has no message to user $n$. In line with the overall massive connectivity scenario, we assume that only a small random subset comprising $K$ out of the $N$ messages are informative, i.e., the frequency of the empty message $\varnothing$ is equal to $\frac{N-K}{N}$. This formulation is reminiscent of the classical broadcast channel, which has been thoroughly studied in information theory. However, the regime that is of interest in massive downlink differs from the widely studied broadcast channel in the following ways:
@@
In the rest of this section, we consider three instances of the massive downlink problem, and demonstrate that, in these fairly general cases, it is possible to significantly reduce the amount of metadata by jointly encoding the messages into a single packet. Specifically, the dependency on $N$ can be reduced from $O(\log N)$ to $O(\log\log N)$ or even $O(1)$ for fairly general problem setups. To keep the presentation clear, we focus on the source coding aspect of the problem and assume that the channel is ideal, i.e., $Y_{n}=X$ for $n=1,\ldots,N$. However, we note that joint encoding also is beneficial from a channel coding perspective, as the larger packet size allows for more efficient coding.
@@ II-D1 Message Acknowledgments
Suppose the base station wants to send binary message acknowledgements to $K$ out of the $N$ users, e.g., after an uplink massive random access transmission. In Fig. 4, this corresponds to the case where $|\mathcal{M}|=1$, so that the base station sends the empty message as a negative acknowledgment and the informative message as a positive acknowledgment. When $K$ is much smaller than $N$, it might seem reasonable to construct the acknowledgment packet by concatenating the identifiers of the users to acknowledge, which would require $K\log_{2}N$ bits. However, as shown in [28], it is possible to significantly better by enumerating all possible subsets of users to acknowledge, which requires only $\log_{2}\binom{N}{K}\leq K\log_{2}(N/K)+K\log_{2}e+O(1)$ bits. Nevertheless, the number of bits can still be significant when $N$ is large. For instance, acknowledging $K=50$ users, each identified by a $64$-bit identifier ($N=2^{64}$), still requires more than $2900$ bits.
@@ II-D2 Transmission of General Feedback Messages
We consider next the problem of transmitting $K$ general messages $W_{1^{\prime}},W_{2^{\prime}},\ldots,W_{K^{\prime}}\in\mathcal{M}$ to $K$ out of the $N$ users, which was recently studied in [29]. Compared to the previous case, instead of assuming a false positive rate, we will assume that the devices know whether their own message is informative or empty, but have no information about the other messages. This assumption is often reasonable, e.g., if the messages are transmitted as responses to an uplink transmission. Alternatively, the message acknowledgment scheme presented in the previous section can be used to indicate the intended recipients. The central problem is then to encode the $K$ messages in such a way that the recipients know how to decode their own message without knowing the identities of the $K-1$ other recipients of informative messages. Note that, if the users know perfectly the identifiers of all $K$ recipients, this can be done simply by listing the messages in the order of the identifiers. However, when each user only knows whether there is a packet for them or not, and not who the other recipients are, the problem becomes more involved. Assuming that the messages are independent and identically distributed across users, the naive encoding scheme of concatenating each of the $K$ messages with the identity of its recipient requires approximately $kH(W_{n})+K\log_{2}N$ bits, where $H(W_{n})$ denotes the entropy of a message $W_{n}$. However as shown in [29], theoretically speaking, the overhead of $K\log_{2}N$ bits can in fact be reduced to only $O(\log_{2}K)$ bits for any independent and identically distributed (i.i.d.) messages, and further to $O(1)$ when the messages are equally likely across the alphabet, thus removing the dependency on $N$.
@@ III-B Enabling Technologies
Finally, THz communications encompass the transmission of signals within the frequency range of $300$ GHz to $3$ THz, offering a significantly increased bandwidth compared to lower frequency bands. This ultra-high bandwidth capacity facilitates the transmission of data at exceptionally high data rates, paving the way for applications that demand both high rate and high reliability, such as wireless virtual reality [39]. However, THz communication encounters limitations, notably its restricted communication range and susceptibility to atmospheric absorption and further research is needed. Although THz communications brings advantages and does have a potential to be combined with RIS, its adoption in 6G standards may still be premature, leaving its use open for future wireless generations.
@@ III-C1 Network Design for Control Systems
  • Age of Information (AoI): Age of Information (AoI) is a metric that measures the time elapsed since the latest relevant information was generated for a state update at the controller. The effectiveness and stability of the control processes hinge on adhering to constraints on AoI statistics, ensuring that the control system operates with the latest and most pertinent state information. The analysis and optimization of AoI statistics has attracted significant attention across diverse communication network scenarios, as evidenced by studies such as [46].
@@
  • Age of Loop (AoL): Age of Loop (AoL) is defined as the elapsed time since the generation of information leading to the latest action or state update in a wireless networked control system (WNCS). Unlike AoI, which is defined for a single communication direction, either downlink or uplink, AoL is considered a more relevant metric for optimizing closed-loop WNCS by considering both uplink and downlink of the control loop [50, 51]. In [50], the AoL metric is employed to learn the WNCS latency and freshness bounds. The study proposes an AoL-based bandwidth allocation policy aimed at minimizing the WNCS cost. Results show that the proposed algorithm outperforms policies based on fixed latency requirements. Expanding on this, [51] conducts a thorough investigation of the peak age of loop (PAoL) performances. This includes the average, variance and outage probability of PAoL for WNCSs operating in the finite blocklength regime over fading channels. This comprehensive analysis paves the way for a PAoL-oriented communication and control co-design, introducing an adaptive scheme for transmission power, blocklength and the maximum number of allowable transmissions. The proposed PAoL-oriented scheme demonstrates significant performance gains when compared to UL-only and DL-only optimization schemes.
@@
  • Value of Information (VoI): Value of Information (VoI) serves as a crucial metric, capturing the variation in a value function concerning the information available to the controller about the state of a process. Distinguished from AoI and AoL, which focus on the timeliness of the updates, VoI delves into the content of the information, measuring the semantics of each data packet [52, 53, 54]. In [52], VoI is quantified as the difference between the benefit derived from enhancing regulation quality due to reduced uncertainty at the controller and the transmission cost of a packet. This study introduces a strategy based on VoI, demonstrating optimal management of the communication between the sensor and controller. It is worth noting that this research assumes negligible network-induced effects, such as transmission delay, quantization or packet dropouts. Extending this concept to networked control systems with arbitrary transmission delay, [53] defines VoI metric as the reduction of uncertainty from the decision maker’s perspective given a measurement update. This VoI metric serves as the triggering condition for NCS closed over a communication network. The proposed VoI based event-triggering policy proves to achieve a lower MSE compared to the AoI-based and periodic updating policies. In the context of an event-triggered control system with multiple subsystems sharing a wireless channel, [54] proposes a channel access prioritization scheme based on the VoI. The VoI of a subsystem depends on channel reliability and represents the difference between sensor’s a posteriori state estimate and the estimator’s estimate in the absence of any data. The objective of the proposed scheme is to minimize the discrepancy in the state estimate. Utilizing measurement based triggering and VoI as the priority measure leads to optimal performance.
@@ III-C2 Network Design for Ultra-Reliable Communication
    In [67], EVT techniques are used at the physical layer to determine the optimum threshold below which the received power samples are classified as extreme events and incorporated into the tail distribution. The generalized Pareto distribution (GPD) is then employed to model this tail distribution. The validity of the proposed EVT-based model is assessed through the examination of probability plots. Building upon this, [68] extends the methodology to characterize the tail of the non-stationary channel distribution. This is achieved by identifying external factors contributing to non-stationarity and segmenting the channel data into multiple stationary sequences. The parameter of the fitting distribution is modeled as a change-point function with respect to time. Furthermore, [69] takes these techniques a step further for systems employing spatial diversity. It calculates the lower tail statistics of received signal power across multiple dimensions, efficiently managing extensive corresponding data. The approach involves fitting bi-variate GPD to the tail of the joint probability distribution, employing logistic distribution-based and Poisson point process-based approaches. The validity of these approaches is evaluated by incorporating the probability measure function of the Pichanks coordinates.
@@ IV-B Slicing of RAN Communication Resources: A Case Study
Figure 6: Typical allocation of time-frequency resources for different kind of traffic typologies in 5G.
@@
For this setup, H-OMA would entail reserving one or more time slots across a number of frequency channels for critical traffic. As pointed out in Section I, though, this may be wasteful in terms of resources because of the sporadic nature of critical-service transmission. The approach in H-NOMA is to let both traffic types overlap in the same time-frequency resources. The idea is to exploit the fact that critical services need to be decoded with a failure rate that is typically $2$ to $3$ orders of magnitude smaller than broadband services. The receiver can exploit this reliability diversity by decoding the critical service first, and then removing the interference it causes on the broadband service prior to its decoding.
@@
Figure 9: A 6G scenario for the coexistence of massive and critical connectivity. A large population of sensors may transmit both ordinary messages and alarm messages to a sink. Alarm messages need to be decoded with much higher reliability than ordinary messages.

### v3

@@ II-C1 Approximate Message Passing Decoding
The choice of the denoiser function in Eq. (4) depends on the specific model assumptions and the goal. Traditionally, the denoiser is implemented as a soft thresholding function, which is suitable when the activation probabilities and $\{g_{n}\}_{n=1}^{N}$ are treated as unknown quantities. However, if the statistics of these quantities are known to the decoder, it is possible to derive minimum mean-square error denoisers, which tend to have better overall performance [16]. The complexity of MMV-AMP iterations is dominated by the computation of $\mathbf{S}\mathbf{X}^{t+1}$, and thus MMV-AMP has a complexity in the order of $O(NLM)$ per iteration [17].
@@ II-C2 Covariance-Based Decoding
The covariance-based approach arises from the observation that the received signal at each antenna are independent and identically distributed realizations of a multivariate complex Gaussian distribution with zero mean and covariance matrix $\mathbf{\Sigma}=\sum_{k=1}^{N}\gamma_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}+\sigma_{Z}^{2}\mathbf{I}$, where the superscript $H$ stands for conjugate transposition. The likelihood function of the received signal $\mathbf{Y}$ given $\boldsymbol{\gamma}=\left[\gamma_{1},\ldots,\gamma_{N}\right]\in\mathbb{R}^{N}$ can then be computed as
$p(\mathbf{Y}|\boldsymbol{\gamma})$ $=\prod_{m=1}^{M}\frac{1}{|\pi\mathbf{\Sigma}|}\exp\left(-\mathbf{y}_{m}^{H}\mathbf{\Sigma}^{-1}\mathbf{y}_{m}\right)$ (6)
@@
Although maximizing the likelihood leads to a non-convex problem, it is possible to obtain good solutions using a coordinate descent algorithm. The resulting set of active users can then be recovered from $\boldsymbol{\gamma}$. By analyzing the coordinate descent, it can be shown that the computational complexity is $O(TNL^{2})$, where $T$ is the number of iterations [25]. In practice, the covariance-based method tends to have longer runtime than AMP for small values of $M$, but is more efficient for larger $M$ [29].
Compared to AMP, performance analysis of the covariance-based approach is more involved, but have been characterized in [29] for the asymptotic case where the number of antennas goes to infinity while $N$, $K$, and $L$ remain finite. In this regime, it is possible to numerically obtain the detection error. This analysis reveals that the algorithm exhibits a phase transition, where the decoding error vanishes for a specific region of $N$, $K$, and $L$. As a key advantage as compared to AMP, the covariance-based approach can work well in the regime where the number of users $K$ is in the order of the square of the preamble length $L$. This is due to the fact that the covariance matrix is of size $O(L^{2})$. However, such advantage comes at a cost of not being able to estimate the channels. Finally, we remark that while the discussion in this section has been confined to the single-cell case, both the algorithmic developments and the analyses of AMP and the covariance method can be extended to the multi-cell or the cell-free massive multiple-input multiple-output (MIMO) scenarios [30, 31, 32, 33].
@@ II-D Massive Downlink Connectivity
In a massive downlink channel, the base station aims to jointly send $N$ messages $W_{1},\ldots,W_{N}$ to users $1,\ldots,N$, respectively, where $W_{n}\in\mathcal{M}\cup\{\varnothing\}$ for $n=1,\ldots,N$. Each message $W_{n}$ ($n=1,\ldots,N$) can either be informative, drawn from the message set $\mathcal{M}$, or empty, denoted by $W_{n}=\varnothing$, indicating that the transmitter has no message to user $n$ or that user $n$ is not actively listening to the base station. In line with the overall massive connectivity scenario, we assume that only a small random subset comprising $K$ out of the $N$ messages are informative, i.e., the fraction of empty messages ($W_{n}=\varnothing$) is equal to $\frac{N-K}{N}$. This formulation is reminiscent of the classical broadcast channel, which has been thoroughly studied in information theory. However, the regime that is of interest in massive downlink differs from the widely studied broadcast channel in the following ways:
@@
In the rest of this section, we consider three instances of the massive downlink problem, and demonstrate that, in these fairly general cases, it is possible to significantly reduce the amount of metadata by jointly encoding the messages into a single packet. Specifically, the dependency on $N$ can be reduced from $O(\log N)$ to $O(\log\log N)$ or even $O(1)$ for fairly general problem setups. To keep the presentation clear, we focus on the source coding aspect of the problem and assume that the channel is ideal, i.e., $Y_{n}=X$ for $n=1,\ldots,N$. However, we note that joint encoding is also beneficial from a channel coding perspective, as the larger packet size allows for more efficient coding.
@@ II-D1 Message Acknowledgments
Suppose the base station wants to send binary message acknowledgements to $K$ out of the $N$ users, e.g., after an uplink massive random access transmission. This corresponds to the case where $|\mathcal{M}|=1$, so that the base station sends the empty message as a negative acknowledgment and the informative message as a positive acknowledgment. When $K$ is much smaller than $N$, it might seem reasonable to construct the acknowledgment packet by concatenating the identifiers of the users to acknowledge, which would require $K\log_{2}N$ bits. However, as shown in [34], it is possible to significantly better by enumerating all possible subsets of users to acknowledge, which requires only $\log_{2}\binom{N}{K}\leq K\log_{2}(N/K)+K\log_{2}e+O(1)$ bits. Nevertheless, the number of bits can still be significant when $N$ is large. For instance, acknowledging $K=50$ users, each identified by a $64$-bit identifier ($N=2^{64}$), still requires more than $2900$ bits.
@@ II-D2 Transmission of General Feedback Messages
We consider next the problem of transmitting $K$ general messages $W_{1^{\prime}},W_{2^{\prime}},\ldots,W_{K^{\prime}}\in\mathcal{M}$ to $K$ out of the $N$ users, which was recently studied in [35, 36]. Compared to the previous case, instead of assuming a false positive rate, we will assume that the devices know whether their own message is informative or empty, but have no information about the other messages. This assumption is often reasonable, e.g., if the messages are transmitted as responses to an uplink transmission. Alternatively, the message acknowledgment scheme presented in the previous section can be used to indicate the intended recipients. The central problem is then to encode the $K$ messages in such a way that the recipients know how to decode their own message without knowing the identities of the $K-1$ other recipients of informative messages. Note that, if the users know perfectly the identifiers of all $K$ recipients, this can be done simply by listing the messages in the order of the identifiers. However, when each user only knows whether there is a packet for them or not, and not who the other recipients are, the problem becomes more involved. Assuming that the messages are independent and identically distributed across users, the naive encoding scheme of concatenating each of the $K$ messages with the identity of its recipient requires approximately $kH(W_{n})+K\log_{2}N$ bits, where $H(W_{n})$ denotes the entropy of a message $W_{n}$. However as shown in [35], theoretically speaking, the overhead of $K\log_{2}N$ bits can in fact be reduced to only $O(\log_{2}K)$ bits for any independent and identically distributed (i.i.d.) messages, and further to $O(1)$ when the messages are equally likely across the alphabet, thus removing the dependency on $N$.
@@ II-D3 Collision-Free Scheduling
II-E Integration of Massive Connectivity in 6G Standardization
So far, we have discussed massive connectivity technologies from a theoretical perspective. In this section, we discuss how these ideas can be integrated in 6G standardization, as well as some of the remaining challenges that need to be addressed.
Let us first focus on the integration of the enhanced technologies for random access schemes in the uplink presented in Section II-B. A first step towards improving the performance would be to reuse the existing access structure in 5G, and augmenting it with, e.g., non-orthogonal preambles. This idea has recently been analyzed for grant-free access in [12], where the authors demonstrate that the inclusion of non-orthogonal preambles in the 5G SDT procedure substantially increases the energy and spectral efficiency, albeit still being far from the theoretical limits. In this case, the sparse recovery algorithms discussed in Section II-C will play a central role in the decoding of the preambles. Furthermore, since grant-based random access also benefits from non-orthogonal pilots, this approach can also be adopted to improve the grant-based random access procedures.
In the context of downlink, the techniques discussed in the previous sections are more straightforward to realize, as they do not require modifications to the physical layer. For the same reason, they do not rely on specific assumptions that need to be validated experimentally. However, while it may be unlikely that downlink designs tailored for specific applications will be implemented in early 6G, the theoretical insights from Section II-D provide useful guidelines in the design of the control information for massive access in 6G, such as the design of acknowledgments and scheduling for grant-based access.
@@ III-B Enabling Technologies
Finally, THz communications encompass the transmission of signals within the frequency range of $100$ GHz to $10$ THz, offering a significantly increased bandwidth compared to lower frequency bands. This ultra-high bandwidth capacity facilitates the transmission of data at exceptionally high data rates, paving the way for applications that demand both high rate and high reliability, such as wireless virtual reality [46]. However, THz communication faces a significant challenge: high atmospheric absorption by atmospheric gases, leading to signal attenuation. Nevertheless, recent advances have addressed this challenge by developing techniques to select frequency bands with lower absorption rates [47], utilize beamforming to avoid regions of high absorption [48], and develop adaptive modulation and coding schemes to minimize errors induced by absorption [49]. These advances have been demonstrated to achieve multi-kilometer transmission ranges [50]. Although THz communications do have a potential to be combined with RIS, its adoption in 6G standards may still be premature, leaving its use open for future wireless generations.
@@ III-C1 Network Design for Control Systems
  • Age of Information (AoI): AoI is a metric that measures the time elapsed since the latest relevant information was generated for a state update at the controller. The effectiveness and stability of the control processes hinge on adhering to constraints on AoI statistics, ensuring that the control system operates with the latest and most pertinent state information. The analysis and optimization of AoI statistics has attracted significant attention across diverse communication network scenarios, as evidenced by studies such as [57].
@@
  • Age of Loop (AoL): AoL is defined as the elapsed time since the generation of information leading to the latest action or state update in a wireless networked control system (WNCS). Unlike AoI, which is defined for a single communication direction, either downlink or uplink, AoL is considered a more relevant metric for optimizing closed-loop WNCS by considering both uplink and downlink of the control loop [61, 62]. In [61], the AoL metric is employed to learn the WNCS latency and freshness bounds. The study proposes an AoL-based bandwidth allocation policy aimed at minimizing the WNCS cost. Results show that the proposed algorithm outperforms policies based on fixed latency requirements. Expanding on this, [62] conducts a thorough investigation of the peak age of loop (PAoL) performances. This includes the average, variance and outage probability of PAoL for WNCSs operating in the finite blocklength regime over fading channels. This comprehensive analysis paves the way for a PAoL-oriented communication and control co-design, introducing an adaptive scheme for transmission power, blocklength and the maximum number of allowable transmissions. The proposed PAoL-oriented scheme demonstrates significant performance gains when compared to UL-only and DL-only optimization schemes.
@@
  • Value of Information (VoI): VoI serves as a crucial metric, capturing the variation in a value function concerning the information available to the controller about the state of a process. Distinguished from AoI and AoL, which focus on the timeliness of the updates, VoI delves into the content of the information, measuring the semantics of each data packet [63, 64, 65]. In [63], VoI is quantified as the difference between the benefit derived from enhancing regulation quality due to reduced uncertainty at the controller and the transmission cost of a packet. This study introduces a strategy based on VoI, demonstrating optimal management of the communication between the sensor and controller. It is worth noting that this research assumes negligible network-induced effects, such as transmission delay, quantization or packet dropouts. Extending this concept to networked control systems with arbitrary transmission delay, [64] defines VoI metric as the reduction of uncertainty from the decision maker’s perspective given a measurement update. This VoI metric serves as the triggering condition for NCS closed over a communication network. The proposed VoI based event-triggering policy proves to achieve a lower MSE compared to the AoI-based and periodic updating policies. In the context of an event-triggered control system with multiple subsystems sharing a wireless channel, [65] proposes a channel access prioritization scheme based on the VoI. The VoI of a subsystem depends on channel reliability and represents the difference between sensor’s a posteriori state estimate and the estimator’s estimate in the absence of any data. The objective of the proposed scheme is to minimize the discrepancy in the state estimate. Utilizing measurement based triggering and VoI as the priority measure leads to optimal performance.
@@ III-C2 Network Design for Ultra-Reliable Communication
    In [78], EVT techniques are used at the physical layer to determine the optimum threshold below which the received power samples are classified as extreme events and incorporated into the tail distribution. The generalized Pareto distribution (GPD) is then employed to model this tail distribution. The validity of the proposed EVT-based model is assessed through the examination of probability plots. Building upon this, [79] extends the methodology to characterize the tail of the non-stationary channel distribution. This is achieved by identifying external factors contributing to non-stationarity and segmenting the channel data into multiple stationary sequences. The parameter of the fitting distribution is modeled as a change-point function with respect to time. Furthermore, [80] takes these techniques a step further for systems employing spatial diversity. It calculates the lower tail statistics of received signal power across multiple dimensions, efficiently managing extensive corresponding data. The approach involves fitting bi-variate GPD to the tail of the joint probability distribution, employing logistic distribution-based and Poisson point process-based approaches. The validity of these approaches is evaluated by incorporating the probability measure function of the Pickands coordinates.
@@ IV-B Slicing of RAN Communication Resources: A Case Study
(d) H-NOMA
Figure 5: (5(a))–(5(c)): Typical allocation of time-frequency resources for different kind of traffic typologies in 5G. (5(d)): Non-orthogonal slicing using H-NOMA.
@@
For this setup, H-OMA would entail reserving one or more time slots across a number of frequency channels for critical traffic. As pointed out in Section I, though, this may be wasteful in terms of resources because of the sporadic nature of critical-service transmission. The approach in H-NOMA is to let both traffic types overlap in the same time-frequency resources, as illustrated in Fig. 5(d). The idea is to exploit the fact that critical services need to be decoded with a failure rate that is typically $2$ to $3$ orders of magnitude smaller than broadband services. The receiver can exploit this reliability diversity by decoding the critical service first, and then removing the interference it causes on the broadband service prior to its decoding.
@@
Figure 8: A 6G scenario for the coexistence of massive and critical connectivity. A large population of sensors may transmit both ordinary messages and alarm messages triggered by a physical phenomenon to a sink. Alarm messages need to be decoded with much higher reliability than ordinary messages.